# Geometric Hardy inequalities on the Heisenberg groups via convexity

### v1

@@ Abstract.
We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$. Our results are based on a geometric condition. This is first implemented for the Euclidean distance in certain non-convex domains. It is also implemented on half-spaces and convex polytopes for the distance defined by the gauge quasi-norm on ${\mathbb{H}}^{n}$ related to the fundamental solution of the horizontal Laplacian. In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains. In all cases the constants are shown to be sharp.
@@
Geometric $L^{p}$-Hardy inequalities; gauge pseudodistance; Heisenberg group; stratified groups of step two; horizontal convexity
@@ 1. Introduction
Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant in the Heisenberg group ${\mathbb{H}}^{n}$ but also on any stratified group of step two. Our approach is based on the general method of [BFT04] and in particular in the $L^{p}$-superharmonicity of the distance function. The general result is then implemented in different contexts.
@@
Another class of distances in a stratified group consists of those induced by any homogenous quasi-norm. In the case of the Heisenberg group $\mathbb{H}^{n}$ one important such quasi-norm is
@@
We recall if we set $Q=2n+2$,the homogeneous dimension of $\mathbb{H}^{n}$, then up to a multiplicative constant, $N^{2-Q}$ is the fundamental solution of the sub-Laplacian on the Heisenberg group ${\mathbb{H}}^{n}$ [Fol73]. Denoting by $d_{N}$ the distance to the boundary incuded by $N$, cf. (12) below, we have the following
@@
The constant is the best possible in both cases.
In Section 4 we extend our setting to that of an arbitrary stratified group of step two. In this setting a certain notion of convexity plays a central role. There are various notions of convexity in the sub-Riemannian setting and their properties can vary significantly [DGN03, DLZ24, LMS03]. Notably, in [MR03] R. Monty and M. Rickly prove that in the case of the Heisenberg group $\mathbb{H}$, if a set is geodesically convex and contains at least three points that do not lie on the same geodesic, then it necessarily coincides with $\mathbb{H}$.
@@
Theorem C. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
Let $\Pi\subset\mathbb{H}^{n}$ be an arbitrary half-space and let $d_{N}(\xi)$, $\xi\in\Pi$, denote the pseudodistance to the boundary with respect to the quasi-norm $N$. Then
$({\rm i})$ The function $d_{N}$ is $C^{1}$ in $\Pi$.
$({\rm ii})$ For any $p>1$ there holds $\Delta_{p,{\mathbb{H}}^{n}}d_{N}\leq 0$ in the distributional sense in $\Pi$.
@@
It is clear from (19) that to prove (i) we only need to restrict out attention near the half-line $\{r=0,\;t>0\}$.
From (15) we easily find that for fixed $t>0$ we have
$s=\frac{t^{\frac{1}{3}}}{r^{\frac{2}{3}}}-\frac{2}{3}\frac{r^{\frac{2}{3}}}{t^{\frac{1}{3}}}+o(r^{\frac{2}{3}}),\qquad\mbox{ as }r\to 0.$
It then follows from (20) that
$G(r,t)=t^{2}-3t^{\frac{4}{3}}r^{\frac{4}{3}}+o(r^{\frac{4}{3}}),\qquad\mbox{ as }r\to 0.$
For the partial derivatives of $G$ we only need to consider the first asymptotic term. From (22) and (23) we arrive at
$G_{r}(r,t)=-4t^{\frac{4}{3}}r^{\frac{1}{3}}+o(r^{\frac{1}{3}})\;,\qquad\quad G_{t}(r,t)=2t+o(1).$
It follows that
$G_{r}(0,t)=0=\lim_{r\to 0}G_{r}(r,t)\;,\qquad G_{t}(0,t)=2t=\lim_{r\to 0}G_{t}(r,t).$
Since the various $o(1)$ above are locally uniform with respect to $t>0$, we conclude that $G$ is $C^{1}$ in $\Pi_{0}$.
To prove (ii) we first note that setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (8),
@@
Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
$A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}_{N}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
@@
We claim that
(27) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
Indeed, suppose to the contary that (27) is not true for some $k$ and some $\xi_{0}\in A_{k}$. Then there exists $\xi_{0}^{\prime}\in\partial\Pi_{k}\setminus E_{k}$ such that
${\rm dist}_{N}(\xi_{0},\partial\Pi_{k})=d_{N}(\xi_{0},\xi_{0}^{\prime}).$
By Lemma 9 there exists a continuous curve joining $\xi_{0}$ and $\xi_{0}^{\prime}$ with the property that each point on the curve has $\xi_{0}^{\prime}$ as its nearest point on $\partial\Pi_{k}$. Since $\xi_{0}^{\prime}\not\in E_{k}$, that curve necessarily intersects some $A_{j}$, $j\neq k$, which is a contradiction.
@@
$d_{N}(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
To prove (i) we fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
$\int_{\Omega}\nabla_{{\mathbb{H}}^{n}}d\cdot\nabla_{{\mathbb{H}}^{n}}\phi\,{\rm d}\xi\geq 0.$
@@
where $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal.
Let $k\in\{1,\ldots,m\}$ be fixed. Integrating by parts and using Proposition 6 we obtain
$\int_{A_{k}}\nabla_{{\mathbb{H}}^{n}}d\cdot\phi\,{\rm d}\xi=$ $-\int_{A_{k}}\phi\,\Delta_{{\mathbb{H}}^{n}}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d\cdot\nu_{H}dS$
$\geq$ $\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d_{k}\cdot\nu_{H}dS.$
We then add over all $k=1,\ldots,m$.
Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$. Let us fix such a set $A_{j}$, $j\neq k$. Denoting by $\nu_{H}$ the horizontal normal vector which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S$ from $A_{k}$ and $A_{j}$ add up to
@@
The surface $S$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have
$(\nabla_{{\mathbb{H}}^{n}}\,d_{k}-\nabla_{{\mathbb{H}}^{n}}\,d_{j})\cdot\nu_{H}$
@@
Combining the above completes the proof of (i). Part (ii) is an immediate consequence of part (i) and Theorem 1. ∎
As an immediate consequence of Theorem 10 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudodistance on ${\mathbb{H}}^{n}$.
@@
Let $D\subset\mathbb{H}^{n}$ be either a bounded convex polytope or an arbitary half-space in ${\mathbb{H}}^{n}$ and let

### v2

@@ Abstract.
We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$. Our results are based on a certain geometric condition. This is first implemented for the Euclidean distance in certain non-convex domains. It is then implemented for the distance defined by the gauge quasi-norm related to the fundamental solution of the horizontal Laplacian when the domain is a half-space or a convex polytope. Finally it is implemented for the Carnot-Carathéodory distance on half-spaces and arbitrary bounded convex domains of ${\mathbb{H}}^{n}$. In all cases the constant $((p-1)/p)^{p}$ is obtained. In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains.
@@
Geometric $L^{p}$-Hardy inequalities; gauge pseudodistance; Carnot-Carathéodory distance; Heisenberg group; stratified groups of step two; horizontal convexity
@@ 1. Introduction
Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant on the Heisenberg group ${\mathbb{H}}^{n}$, but also on any stratified group of step two. Our approach is based on the general method of [BFT04] and in particular in the $L^{p}$-superharmonicity of the distance function. The general result is then implemented in different contexts.
@@
Another class of distances in a stratified group consists of those induced by a homogenous quasi-norm. In the case of the Heisenberg group $\mathbb{H}^{n}$ there are two particularly important such (quasi)-norms. The first is
@@
We note that $N^{2-Q}$, with $Q=2n+2$ being the homogeneous dimension of $\mathbb{H}^{n}$, is (up to a multiplicative constant) the fundamental solution of the sub-Laplacian on ${\mathbb{H}}^{n}$ [Fol73].
A second important homogeneous quasi-norm, which respects the sub-Riemannian geometry of ${\mathbb{H}}^{n}$, is the one arising from the Carnot-Carathéodory distance, denoted here by $\rho$. Let us recall the definition of the Carnot-Carathéodory distance between two points $x,y\in\mathbb{R}^{n}$: For a family $\{X_{1},\cdots,X_{m}\}$ of vector fields, the Carnot-Carathéodory distance between $x,y$ is given by
(2) $\rho(x,y)=\inf_{\gamma\in\mathcal{C}_{x,y}}\{{\rm length}(\gamma)\}\,,$
where $\mathcal{C}_{x,y}$ is the set of horizontal curves joining $x$ to $y$. In the case of a stratified group when the set $\{X_{1},\cdots,X_{m}\}$ coincides with the first stratum $V_{1}$ of its Lie algebra (see Section 4 for related definitions) we always have $d(x,y)<\infty$, and the Carnot-Carathéodory distances induced by different choices of $V_{1}$ are equivalent [Pa89, Section 1.3].
Denoting by $d_{N}$ the distance to the boundary induced by $N$, cf. (13) below, we have the following
@@
Moreover the constant is the best possible in both cases.
Denoting by $d_{\rho}$ the Carnot-Carathéodory distance to the boundary, we obtain the following
Theorem C. Let $p>1$ and let $D\subset{\mathbb{H}}^{n}$ be a bounded and convex domain or a half-space in ${\mathbb{H}}^{n}$. There holds
$\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\left(\frac{p-1}{p}\right)^{p}\int_{D}\frac{|u|^{p}}{d_{\rho}^{p}}\,{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(D)\,.$
In the case of the half-space the constant is sharp.
The appereance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e. A more interesting difference between the above two homogeneous norms relates to nearest boundary points in the case of a subspace $\Pi\subset{\mathbb{H}}^{n}$. The correspondence
$\Pi\ni\xi\longleftrightarrow(\xi^{\prime},s)\in\partial\Pi\times\mathbb{R}_{+}$
where $s={\rm dist}(\xi,\partial\Pi)=d(\xi,\xi^{\prime})$ is different in the two cases. In the case of the quasi-norm (1) the above is a simple 1-1 correspondance, as in the Euclidean case, cf. Proposition 9. However in the case of the Carnot-Carathéodory distance the situation is entirely different; see Proposition 16.
In Section 4 we extend our setting to that of an arbitrary stratified group of step two. In this general setting, a certain notion of convexity plays a central role. There are various notions of convexity in the sub-Riemannian setting and their properties can vary significantly [DGN03, DLZ24, LMS03]. Notably, in [MR03] R. Monty and M. Rickly prove that in the case of the Heisenberg group $\mathbb{H}$, if a set is geodesically convex and contains at least three points that do not lie on the same geodesic, then it necessarily coincides with $\mathbb{H}$.
@@
Theorem D. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
Let $\Pi\subset\mathbb{H}^{n}$ be an arbitrary half-space and let $d_{N}(\xi)$, $\xi\in\Pi$, denote the pseudodistance to the boundary with respect to the quasi-norm $N$. Then for any $p>1$ there holds $\Delta_{p,{\mathbb{H}}^{n}}d_{N}\leq 0$ in the distributional sense in $\Pi$.
@@
Setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (9),
@@
For simplicity we write $d$ instead of $d_{N}$. Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
$A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
@@
We then have
(28) $d(\xi)={\rm dist}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
@@
$d(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
We fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
(29) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
@@
where
(30) $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$
and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal vector.
We cannot directly apply integration by part to each of the integrals in the RHS of (29) due to the fact that, as is seen from Lemma 5, there is a halfline $L_{k}\subset\Pi_{k}$ on which each function $d_{k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$). We define the set
$K=\bigcup_{k=1}^{m}L_{k}$
and so $d$ is $C^{1}$ in $\Omega\setminus K$.
We now use a standard cut-off argument. We denote by $d_{E}(\cdot,K)$ the Euclidean distance to the set $K$ and we consider smooth functions $\psi_{\epsilon}$, $\epsilon>0$, such that $0\leq\psi_{\epsilon}\leq 1$ and
$\left\{\begin{array}[]{ll}\psi_{\epsilon}(\xi)=0,&\mbox{ if }d_{E}(\xi,K)<\epsilon,\\[2.84544pt] \psi_{\epsilon}(\xi)=1,&\mbox{ if }d_{E}(\xi,K)>2\epsilon,\\[2.84544pt] |\nabla\psi_{\epsilon}(\xi)|\leq\frac{c}{\epsilon},&\mbox{ for all }\xi\in{\mathbb{H}}.\end{array}\right.$
We then define $\phi_{\epsilon}=\phi\,\psi_{\epsilon}$.
Let $k\in\{1,\ldots,m\}$. Integrating by parts and using Proposition 6 we obtain
$\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS$
$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
where $\nu_{k,H}$ is defined as above relative to $A_{k}$. Adding over $k$ we arrive at
(31) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$. Let us fix $k,j$ such a set $A_{k}$ and $A_{j}$ share such a common boundary $S_{kj}$. Letting, as above, $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ where $\nu$ is the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
@@
The surface $S_{kj}$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have on $S_{kj}$
$(\nabla_{H}\,d_{k}-\nabla_{H}\,d_{j})\cdot\nu_{H}$
@@
Hence the LHS of (31) is non-negative. Noting that $\nabla\phi_{\epsilon}\to\nabla\phi$ in $L^{1}(\Omega)$ as $\epsilon\to 0$ completes the proof. Combining the above completes the proof of (i). Part (ii) is an immediate consequence of part (i) and Theorem 1.The optimality of the constant follows from Theorem 8.∎
As an immediate consequence of Theorem 10 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudo-distance on ${\mathbb{H}}^{n}$.
@@
Let $D\subset\mathbb{H}^{n}$ be either a bounded convex polytope or an arbitrary half-space in ${\mathbb{H}}^{n}$ and let